OpenStudy (anonymous):

We have a bug moving alone the curve y = 4 - x^2/16. The distance is measured in feet. When the bug passes through point (4,3) its y-coordinate is increasing at 20 ft/sec so at what rate is its x-coordinate changing?

OpenStudy (anonymous):

:/ sorry not good with this kind of stuff, i could call for help tho @taylorannbs @makayla_finkbeiner_ @Bookworm14 @aprilann143

OpenStudy (anonymous):

wow thanks man

OpenStudy (anonymous):

no prob

ganeshie8 (ganeshie8):

you're given : \(\dfrac{dy}{dt} = 20\)

ganeshie8 (ganeshie8):

you need to find \(\dfrac{dx}{dt}\)

ganeshie8 (ganeshie8):

differentiate the given function with.respect.to \(t\)

OpenStudy (anonymous):

wait this is lagging

ganeshie8 (ganeshie8):

yeah, just differentiate ur function, plugin given values, solve for dx/dt

OpenStudy (anonymous):

so that is t/x + 1/8t^2 with implicit diff right?

ganeshie8 (ganeshie8):

\(y = 4 - \dfrac{x^2}{16}\)

ganeshie8 (ganeshie8):

Is that ur given curve ?

OpenStudy (anonymous):

ok so i need to find dx/dt of that right?

OpenStudy (anonymous):

I believe so, that would give me -t^2(x) right?

ganeshie8 (ganeshie8):

what do u mean u believe so ? you're not given the question/equation in correct format ?

ganeshie8 (ganeshie8):

whats wid two `//` in the question ? `y = 4 - x^2//16` what kind of operator is a `//` ?

OpenStudy (anonymous):

ohhhh that was an accident

ganeshie8 (ganeshie8):

lol okay, take a snapshot of the question and attach if psble

OpenStudy (anonymous):

i changed it and made it right

ganeshie8 (ganeshie8):

okay :)

ganeshie8 (ganeshie8):

\(y = 4 - \dfrac{x^2}{16}\) differentiate both sides with.respect.to \(t\) : \(\dfrac{dy}{dt} = 0 - \dfrac{1}{16}(2x)\dfrac{dx}{dt} \)

ganeshie8 (ganeshie8):

^^simplify

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

lol i'm horrible with this but I got x/2t + 4t/x

OpenStudy (anonymous):

wait no dy/dt = -32x dx/dt

ganeshie8 (ganeshie8):

could u show me how u got that ?

ganeshie8 (ganeshie8):

careful, 2 is in numerator and 16 is in denominator

ganeshie8 (ganeshie8):

\(\dfrac{dy}{dt} = 0 - \dfrac{1}{16}(2x)\dfrac{dx}{dt} \) simplifies to \(\dfrac{dy}{dt} = - \dfrac{x}{8}\dfrac{dx}{dt} \)

OpenStudy (anonymous):

oh well i got the other one by simplifying 0-1/16(2x)

ganeshie8 (ganeshie8):

plugin the given values : \(x = 4\) \(\dfrac{dy}{dt} = 20\) and solve \(\dfrac{dx}{dt}\)

ganeshie8 (ganeshie8):

\(\dfrac{dy}{dt} = - \dfrac{x}{8}\dfrac{dx}{dt} \) \(20 = - \dfrac{4}{8}\dfrac{dx}{dt} \)

ganeshie8 (ganeshie8):

solve \(\dfrac{dx}{dt}\)

OpenStudy (anonymous):

it's not -40 is it?

ganeshie8 (ganeshie8):

-40 is right !

OpenStudy (anonymous):

no way

OpenStudy (anonymous):

that's final?

ganeshie8 (ganeshie8):

thats final.

OpenStudy (anonymous):

well thanks a lot i mean i got a study guide and i need the help to get ready for this final